OpenStudy (keenan):

4x+3>=2 need two answers

OpenStudy (keenan):

thats one

OpenStudy (anonymous):

in absolute value he wants the absolute value and the negative reciprocal for like graphing

OpenStudy (sriram):

do u mean absolute value??

OpenStudy (keenan):

yes

OpenStudy (anonymous):

like flip the problem and signs

OpenStudy (keenan):

do you have answers

OpenStudy (sriram):

wait i think i have made some mistake

OpenStudy (sriram):

i got the mistake

OpenStudy (sriram):

mod(4x+3)>=2 hence when 4x-3>0 the mod would be removed to give 4x+3>2 and X> -1/4 but this is only when 4x+3>0 that is for x lying btwn (-3/4 , infinity) we now take the intersection of the 2 results we get the intersection as (-1/4, infinity) thats one

OpenStudy (sriram):

the other one would be when when 4x+3<0 ie when x lies (-infinity , -3/4) at that moment on removing the mod we get -4x-3>=2 hence x = (-infinity, -5/4) taking intersection (-infinity,-5/4)

OpenStudy (sriram):

sorry forgot the equal to sign hence x= ( -infinity , -5/4] U [-1/4, infinity)

OpenStudy (anonymous):

\[|4x+3|\geq 2\] solve \[4x+3\geq 2\] and \[4x+3\leq -2\] separately that is all

OpenStudy (anonymous):

think about what you are actually analyzing in an absolute value inequality, this will allow you to approach these problems more insightfully.